Write an equation in standard form using only integers for the line described .
The line perpendicular to y = 3/7x +7 and containing (3,-7)

Answer by ewatrrr(24785)

P(3,-7), m = -7/3
y + 7 = (-7/3)(x - 3)
y = (-7/3)x + 7 - 7
y = (-7/3)x
7x + 3y = 0